# In-Context Reverse Classification Accuracy: Efficient Estimation of Segmentation Quality without Ground-Truth

### v1

@@ Abstract
Assessing the quality of automatic image segmentation is crucial in clinical practice, but often very challenging due to the limited availability of ground truth annotations. In this paper, we introduce In-Context Reverse Classification Accuracy (In-Context RCA), a novel framework for automatically estimating segmentation quality in the absence of ground-truth annotations. By leveraging recent in-context learning segmentation models and incorporating retrieval-augmentation techniques to select the most relevant reference images, our approach enables efficient quality estimation with minimal reference data. Validated across diverse medical imaging modalities, our method demonstrates robust performance and computational efficiency, offering a promising solution for automated quality control in clinical workflows, where fast and reliable segmentation assessment is essential. The code is available at https://github.com/mcosarinsky/In-Context-RCA.
1 Introduction
@@
Once a segmentation method is deployed in practice, continuously assessing its real performance on new unseen data samples becomes a challenging task, mainly due to the limited availability of reference annotations in clinical settings. However, evaluation remains a critical task, especially for identifying instances where an image segmentation method may fail in automatic pipelines due to unexpected changes, e.g. in patient demographics[10] or acquisition protocols. Here we aim to design a method that can quickly and automatically assess the quality of segmentations when there is no GT available to compare with.
@@
We propose a flexible framework based on In-Context segmentation [32] and Reverse Classification Accuracy (RCA) [31] that we name In-Context RCA. RCA is a framework originally designed to predict the performance of an image segmentation method in scenarios where GT masks are unavailable. It works by using the predicted segmentation of a new image as a pseudo-ground truth to train a reverse classifier (a segmentation network in our case) [^1] Note that we use the term reverse classifier as introduced in the original publication [31], although it would be more accurately described as a reverse segmenter. We will use both terms interchangeably.. This model is then applied to a set of reference images with known GT, and its performance is evaluated using standard metrics like the DSC. Here we propose to combine this framework with In-Context Learning (ICL), as it naturally handles such few-shot scenarios, proposing to eliminate the need of tranining a reverse segmenter model. ICL is a paradigm that allows models to adapt to new tasks by conditioning on a small set of labeled examples, without requiring additional training. For image segmentation, it means that we can segment a target image by providing a set of example images with their corresponding segmentation masks at inference time, without re-training our segmenter [4, 37]. By leveraging in-context segmentation models as reverse classifiers, our method facilitates segmentation quality estimation at the individual image level.
@@
Related work. Despite its relevance, few approaches have addressed automatic quality control of segmentations thus far. Many previous works have relied on extracting features from segmentations and fitting a model to these features. Such approaches have been applied to prostate segmentation [29], cardiac segmentation [33], and a more generic framework has been used to assess the quality for automated segmentations for lung and liver in [13]. However, these methods are constrained by the feature extraction, which can limit their adaptability and generalizability to diverse clinical data. Additionally, a CNN-based approach was developed for automated quality control of left ventricle segmentation in cardiac MRI [9]. In spite of demonstrating the potential of deep learning for automatic quality assessment, this method requires generating a synthetic dataset and further training multiple CNN models.
@@
Other previous studies rely on a reverse-testing strategy, where a new classifier is trained on predictions on the test data and evaluated again on the training data [8, 36]. The limitation of these approaches is that they only provide an estimate of the average performance rather than case-specific estimations. Building on this idea, Reverse Classification Accuracy (RCA) has been shown to provide accurate predictions of segmentation quality on a per-image basis [31, 24]. In [31], Atlas-based segmentation has shown to be the most effective reverse classifier for medical image segmentation. However, the computational time required to complete atlas based segmentation on a single instance is a limitation, making it impractical for real-time quality control. In [23] the authors propose using a neural network to directly estimate prediction quality, aiming to mitigate the time constraint. However, this approach necessitates training a separate CNN for each task, which significantly limits its practicality and scalability. More importantly, both methods [31, 24] have only been used for anatomical segmentation and not for lesions, which are usually more challenging due to the lack of regularity.
@@
Recently, Few-shot Semantic Segmentation (FSS) methods –which extend in-context learning to address segmentation tasks– have shown great potential in the medical domain, particularly with limited annotations [19]. While traditional approaches are often based on prototypical networks [27, 18], more advanced models such as UniverSeg [4] and foundational models like SAM 2 [22, 3, 37] delivered more promising results on a wide range of medical tasks. This enhanced generalizability is largely due to their training on vast amounts of diverse data. Here we will use FSS to improve RCA performance and efficiency.
Contributions. We propose a novel approach titled In-Context RCA, which extends the traditional RCA framework by leveraging in-context segmentation models as reverse classifiers. These models can adapt their predictions based on a limited number of labeled examples without the need for additional training, making them ideal for this setting. Furthermore, inspired by the recent success of retrieval augmented generation frameworks [14] for natural language processing, we introduce a more efficient variant of RCA entitled Retrieval Augmented RCA, which incorporates image embedding vector databases to efficiently retrieve the most relevant reference images to be used as few-shot examples, further enhancing efficiency and accuracy. We provide extensive validation using 10 medical imaging datasets, covering a wide range of image modalities and organs, broadening the scope of RCA analysis conducted so far and demonstrating the method’s capability to provide accurate estimates in a wide range of scenarios while drastically reducing the required inference time.
2 In-Context RCA
@@
Let us say we deployed to production an X-ray lung segmentation model, and we want to monitor the quality of the segmentations $S_{I}$ produced for new images $I$. Let us also assume that we have at our disposal a database of reference images $J_{i}$ and GT segmentation masks $S_{J_{i}}^{GT}$ (in our case, X-ray images with expert annotations). The core idea behind RCA is to assess the quality of a segmentation $S_{I}$ corresponding to a new image $I$ by training a reverse segmenter $f_{I,S_{I}}$ (or reverse classifier) using only $S_{I}$ as pseudo-ground truth for training. The reverse segmenter is then applied to the reference database containing images $J_{i}$ with known GT masks $S_{J_{i}}^{GT}$, generating segmentations $f_{I,S_{I}}(J_{i})=S_{J_{i}}$. The quality of these segmentations is then assessed by computing a chosen quality metric $\rho$ (e.g DSC) for each $S_{J_{i}}$ and its corresponding ground truth $S_{J_{i}}^{GT}$. Finally, we estimate the metric $\rho$ for $S_{I}$ as the best metric value on the reference database:
@@
Figure 1: General outline of the In-Context RCA framework. The quality of a predicted segmentation $S_{I}$ on an image $I$ is estimated by leveraging $(I,S_{I})$ as the support set of an In-Context Classifier or segmenter. This classifier is then applied to segment a reference dataset, selected through retrieval augmentation. Finally, the best segmentation result based on an evaluation metric $\rho$, is used to predict the quality of $S_{I}$ following Eq 1.
@@
Here we adopt UniverSeg as the reverse classifier of the In-Context RCA framework. To this end, we use the target image $I$ and the predicted segmentation under assessment $S_{I}$ as the UniverSeg support set, i.e. we simply define $\mathcal{S}=\{(I,S_{I})\}$. We then segment all the reference images $J_{i}$ in our RCA reference database by considering them as query images, one at a time, and finally compare each resulting segmentation $S_{J_{k}}$ with the corresponding reference segmentation $S_{J_{k}}^{GT}$ to estimate $\hat{\rho}(S_{I})$ as defined in Eq. 1.
SAM 2-based reverse segmenter. Segment Anything Model (SAM) [12] is a foundation model for image and video segmentation guided by prompts. Its architecture consists of image and prompt encoders that extract features, and a mask decoder that fuses this information to generate segmentations. SAM 2 [22] extends its predecessor by incorporating memory components (an encoder, bank, and attention mechanism) enabling adaptation to new tasks and outperforming existing video object segmentation models, especially for temporal tracking. For medical FSS [3], SAM 2 treats the support set as initial frames of a sequence, with their segmentation masks serving as prompts. Despite having no temporal connection between images, this approach effectively transfers segmentation masks from annotated to unannotated examples (see [3] for more details). In our framework, we use SAM 2 as reverse classifier, following the few-shot setting with sequences that comprise only two frames: 1) the target image and its pseudo-ground truth, and 2) a query image drawn from the reference dataset.
Retrieval Augmented (RA)-RCA. We propose to reduce the size of the RCA reference database by dynamically retrieving the most relevant reference images on a per-case basis. This allows the model to focus on smaller, more relevant subsets of data, optimizing computational efficiency while maintaining and even improving the quality of predictions. To implement this, we embed all the images from the reference database using DINOv2 embeddings [17] and then create a vector database using FAISS [7], which enables efficient similarity search. For each incoming target image whose segmentation quality we want to assess, we compute its embedding and retrieve the $k$ most similar images from the reference index based on cosine similarity. While DINOv2 was not explicitly trained for the medical domain, it has shown strong generalization, making it effective for general tasks, including those in medical imaging. In fact, a similar retrieval-augmented approach for few-shot medical image segmentation has been explored in [34], demonstrating its effectiveness in similar scenarios. In the case of X-ray images, we leverage RAD-DINO [20], a domain-specific version of DINOv2 fine-tuned for radiology, as we expect it to further improve performance.
3 Experiments and results
Datasets. We evaluate In-Context RCA across diverse datasets listed in Figure 2 including SCD [21], HC18 [11], PSFHS [2], PH2 [15], ISIC 2018 [5], NuCLS [1], JSRT [26], 3D-IRCAdB [28], WBC-CV and WBC-JTSC [35]. Unlike previous studies that explored RCA focusing on cardiac MRI [24, 23] or multi-organ segmentation on whole-body MRI [31] only, we broaden our analysis to a wider range of medical imaging challenges. To maintain consistency, we convert all images to grayscale and resize them to 256x256 pixels. For 3D images, we extract slices, keeping those where the region of interest (ROI) is present.

### v3

# Conformal In-Context Reverse Classification Accuracy: Efficient Estimation of Segmentation Quality with Statistical Guarantees

@@ Abstract
Assessing the quality of automatic image segmentation is crucial in clinical practice, but often very challenging due to the limited availability of ground truth annotations. Reverse Classification Accuracy (RCA) is an approach that estimates the quality of new predictions on unseen samples by training a segmenter on those predictions, and then evaluating it against existing annotated images. In this work, we introduce Conformal In-Context RCA, a novel method for automatically estimating segmentation quality with statistical guarantees in the absence of ground-truth annotations, which consists of two main innovations. First, In-Context RCA, which leverages recent in-context learning models for image segmentation and incorporates retrieval-augmentation techniques to select the most relevant reference images. This approach enables efficient quality estimation with minimal reference data while avoiding the need of training additional models. Second, we introduce Conformal RCA, which extends both the original RCA framework and In-Context RCA to go beyond point estimation. Using tools from split conformal prediction, Conformal RCA produces prediction intervals for segmentation quality providing statistical guarantees that the true score lies within the estimated interval with a user-specified probability. Validated across 10 different medical imaging tasks in various organs and modalities, our methods demonstrate robust performance and computational efficiency, offering a promising solution for automated quality control in clinical workflows, where fast and reliable segmentation assessment is essential. The code is available at https://github.com/mcosarinsky/Conformal-In-Context-RCA.
I Introduction
@@
Once a segmentation method is deployed in practice, continuously assessing its real performance on new unseen data samples becomes a challenging task, mainly due to the limited availability of reference annotations in clinical settings. However, evaluation remains a critical task, especially for identifying instances where an image segmentation method may fail in automatic pipelines due to unexpected changes, e.g. in patient demographics[4] or acquisition protocols. This motivates the development of methods that can automatically estimate segmentation quality when no ground truth is available.
@@
Reverse Classification Accuracy (RCA) [5] is a framework originally designed to predict the performance of an image segmentation method in scenarios where GT masks are unavailable. It works by using the predicted segmentation of a new image as a pseudo-ground truth to train a reverse classifier (a segmentation network in our case) [^1] Note that we use the term reverse classifier as introduced in the original publication [5], although it would be more accurately described as a reverse segmenter. We will use both terms interchangeably.. This model is then applied to a set of reference images with known GT, and its performance is evaluated using standard metrics like the DSC. The distribution of scores over the reference set serves as a proxy for the quality of the original segmentation. Even though RCA has been widely adopted in the medical imaging community, it still suffers from two main limitations: it is very inefficient in computational terms and it lacks a mechanism to understand how reliable the estimated segmentation performance is. In this work, we focus on addressing both limitations by proposing a flexible approach that combines recent advances in In-Context Learning (ICL) for segmentation [6], with split conformal prediction [7, 8] to build confidence intervals for the estimated metrics.
We first introduce In-Context RCA, which leverages in-context learning segmentation models [9, 10] that can perform new tasks by conditioning on a small set of labeled examples without requiring any additional training. In the context of image segmentation, this means we can segment a new image by simply providing a few example image-mask pairs, avoiding the need to train a new segmentation model. By leveraging in-context segmentation models as reverse classifiers, our framework makes the process more efficient and practical for clinical use.
More importantly, we address a key limitation that is shared across most existing segmentation quality control frameworks: the lack of statistical guarantees for the estimated quality scores. Traditionally RCA yields a single point estimate of segmentation quality, such as the mean or maximum over the set of reference scores. While useful, this estimate does not express the uncertainty inherent in the prediction. In this work, we propose to generate confidence intervals by incorporating tools from split conformal prediction [7, 8, 11], which provide statistical guarantees that the intervals contain the true segmentation score with a user-specified probability. This results in a more informative and reliable assessment of segmentation quality, going beyond point estimates to quantify uncertainty at the individual image level.
II Related work
II-A InContext Learning and Reverse Classification Accuracy
@@
Despite its relevance, few approaches have addressed automatic quality control of segmentations thus far. Many previous works have relied on extracting features from segmentations and fitting a model to these features. Such approaches have been applied to prostate segmentation [12], cardiac segmentation [13], and a more generic framework has been used to assess the quality for automated segmentations for lung and liver in [14]. However, these methods are constrained by the feature extraction, which can limit their adaptability and generalizability to diverse clinical data. Additionally, a CNN-based approach was developed for automated quality control of left ventricle segmentation in cardiac MRI [15]. In spite of demonstrating the potential of deep learning for automatic quality assessment, this method requires generating a synthetic dataset and further training multiple CNN models.
@@
Other previous studies rely on a reverse-testing strategy, where a new classifier is trained on predictions on the test data and evaluated again on the training data [16, 17, 18]. The limitation of these approaches is that they only provide an estimate of the average performance rather than case-specific estimations. Building on this idea, Reverse Classification Accuracy (RCA) has been shown to provide accurate predictions of segmentation quality on a per-image basis [5, 19]. In [5], Atlas-based segmentation has shown to be the most effective reverse classifier for medical image segmentation. However, the computational time required to complete atlas based segmentation on a single instance is a limitation, making it impractical for real-time quality control. In [20] the authors propose using a neural network to directly estimate prediction quality, aiming to mitigate the time constraint. However, this approach necessitates training a separate CNN for each task, which significantly limits its practicality and scalability. More importantly, both methods [5, 19] have only been used for anatomical segmentation and not for lesions, which are usually more challenging due to the lack of regularity.
@@
Recently, Few-shot Semantic Segmentation (FSS) methods —which extend in-context learning to address segmentation tasks— have shown great potential in the medical domain, particularly with limited annotations [21]. While traditional approaches are often based on prototypical networks [22, 23], more advanced models such as UniverSeg [9] and foundational models like SAM 2 [24, 25, 10] delivered more promising results on a wide range of medical tasks. This enhanced generalizability is largely due to their training on vast amounts of diverse data. Here we will use FSS to improve RCA performance and efficiency.
II-B Performance Interval Estimation
While most existing methods for segmentation quality estimation focus on predicting point estimates, few works aim to quantify the uncertainty of such estimations. Recently, the idea of estimating performance ranges has emerged as a promising direction. One notable approach is the method proposed in [26], which applies split conformal prediction to produce intervals for segmentation quality scores with formal coverage guarantees. To do so, however, they must either train specialised probabilistic segmentation models that can sample multiple plausible outputs for the same input—such as Probabilistic U‑Net [27] or PHiSeg [28]—or rely on alternative uncertainty estimation techniques like test‑time augmentation [29], Monte‑Carlo dropout [30], or ensembling [31]; all of which add further implementation overhead and yielded lower performance in the reported experiments. In addition, the method constructs its intervals from nonconformity scores based on the softmax scores [32, 33], which are known to suffer from calibration issues [34, 35]. In this work, we introduce Conformal RCA, which is model-agnostic and can be used to estimate quality intervals for segmentation masks regardless of the underlying architecture or training procedure used to produce them. Unlike [26], we base our conformal prediction framework on Conformalized Quantile Regression (CQR), leveraging its ability to produce tighter prediction intervals than alternative conformal approaches commonly used for regression [11], and introducing novel modifications tailored to the RCA procedure.
Our Contributions:
  • We propose In-Context RCA, extending the traditional RCA framework by leveraging in-context segmentation models as reverse classifiers. These models adapt their predictions using a few labeled examples at inference time, without requiring additional training.
  • We introduce Retrieval-Augmented RCA, a more efficient variant of RCA that uses embedding-based retrieval to select the most relevant reference images to be used as few-shot examples, enhancing efficiency and accuracy.
  • We propose to incorporate conformal prediction into RCA for the first time, enabling the estimation of prediction intervals with statistical guarantees, and providing image-level uncertainty quantification. We base our approach on CQR, but unlike prior work that trains a quantile regression model, we estimate the quantiles directly from the RCA reference set, avoiding additional training.
  • We conduct extensive experiments on 10 diverse medical imaging datasets, demonstrating strong performance across different modalities and anatomy. We significantly broaden the scope of RCA while reducing inference time.
Figure 1: General outline of the In-Context RCA framework. The quality of a predicted segmentation $S_{I}$ on an image II is estimated by leveraging $(I,S_{I})$ as the support set of an In-Context Classifier or segmenter. This classifier is then applied to segment a reference dataset, selected through retrieval augmentation. Finally, the best segmentation result based on an evaluation metric $\rho$, is used to predict the quality of $S_{I}$ following Eq. (2).
III In-Context RCA
@@
Suppose we deployed to production an X-ray lung segmentation model, and we want to monitor the quality of the segmentation $S_{I}$ produced for a new image $I$. Let us also assume access to a reference database $D=\{(J_{i},S_{J_{i}}^{GT})\}_{i=1}^{m}$ of annotated images (in our case, X-ray images with expert annotations). The core idea behind RCA is to assess the quality of a segmentation $S_{I}$ corresponding to a new image $I$ by training a reverse segmenter $f_{I,S_{I}}$ (or reverse classifier) using only $S_{I}$ as pseudo-ground truth for training. The reverse segmenter is then applied to each image in $D$, generating segmentations $f_{I,S_{I}}(J_{i})=S_{J_{i}}$. The quality of these segmentations is then assessed by computing a chosen quality metric $\rho$ (e.g DSC) for each $S_{J_{i}}$ and its corresponding ground truth $S_{J_{i}}^{GT}$. This process yields a set of scores:
$\mathcal{Y}=\left\{y^{(k)}(S_{I})=\rho(S_{J_{k}},S_{J_{k}}^{GT})\right\}_{k=1}^{m}.$ (1)
Traditionally, RCA collapses this set into a point estimate of segmentation quality by taking the best metric value (or the mean) over $\mathcal{Y}$:
@@
Here we adopt UniverSeg as the reverse classifier of the In-Context RCA framework. To this end, we use the target image $I$ and the predicted segmentation under assessment $S_{I}$ as the UniverSeg support set, i.e. we simply define $\mathcal{S}=\{(I,S_{I})\}$. We then segment all the reference images $J_{i}$ in our RCA reference database by considering them as query images, one at a time, and finally compare each resulting segmentation $S_{J_{k}}$ with the corresponding reference segmentation $S_{J_{k}}^{GT}$ to estimate $\hat{\rho}(S_{I})$ as defined in Eq. (2).
SAM 2-based reverse segmenter. Segment Anything Model (SAM) [36] is a foundation model for image and video segmentation guided by prompts. Its architecture consists of image and prompt encoders that extract features, and a mask decoder that fuses this information to generate segmentations. SAM 2 [24] extends its predecessor by incorporating memory components (an encoder, bank, and attention mechanism) enabling adaptation to new tasks and outperforming existing video object segmentation models, especially for temporal tracking. For medical FSS [25], SAM 2 treats the support set as initial frames of a sequence, with their segmentation masks serving as prompts. Despite having no temporal connection between images, this approach effectively transfers segmentation masks from annotated to unannotated examples (see [25] for more details). In our framework, we use SAM 2 as a reverse classifier, following the few-shot setting with sequences that comprise only two frames: 1) the target image and its pseudo-ground truth, and 2) a query image drawn from the reference dataset.
@@
III-A Retrieval Augmented (RA)-RCA
In order to further improve the efficiency of RCA, we propose to reduce the size of the RCA reference database by dynamically retrieving the most relevant reference images on a per-case basis. This allows the model to focus on smaller, more relevant subsets of data, optimizing computational efficiency while maintaining and even improving the quality of predictions. To implement this, we embed all the images from the reference database using DINOv2 embeddings [37] and then create a vector database using FAISS [38], which enables efficient similarity search. For each incoming target image whose segmentation quality we want to assess, we compute its embedding and retrieve the $k$ most similar images from the reference index based on cosine similarity. While DINOv2 was not explicitly trained for the medical domain, it has shown strong generalization, making it effective for general tasks, including those in medical imaging. In fact, a similar retrieval-augmented approach for few-shot medical image segmentation has been explored in [39], demonstrating its effectiveness in similar scenarios. In the case of X-ray images, we leverage RAD-DINO [40], a domain-specific version of DINOv2 fine-tuned for radiology, as we expect it to further improve performance.
IV Conformal RCA
IV-A Preliminaries: Split Conformal Prediction
Conformal Prediction (CP) is a model-agnostic, distribution-free framework for uncertainty quantification that provides statistically valid prediction sets or intervals, for classification and regression tasks respectively, under minimal assumptions [7, 8]. Given a desired coverage level $1-\alpha$ with $\alpha\in(0,1)$ and a “nonconformity score” function, CP constructs a prediction region $C_{\alpha}(X)$ for a given sample $X$ such that $Y\in C_{\alpha}(X)$ with probability $1-\alpha$, where $Y$ is the true label. Here, the desired coverage level $1-\alpha$ specifies the proportion of times the true outcome is expected to fall within the prediction set, while the nonconformity score quantifies how unusual or “nonconforming” a candidate prediction is with respect to the observed data.
We focus on split conformal prediction, a practical variant that conformalizes a fixed model using a held-out calibration set. This contrasts with full conformal prediction, which typically requires retraining the model for each new test point. The only required assumption is that the calibration samples and the test samples are exchangeable (e.g., the samples are i.i.d.). Let us now state the general procedure for split conformal prediction, as presented in [8]. Given a calibration set $\{(x_{i},y_{i})\}_{i=1}^{n}$ and a new test input $x_{n+1}$:
  1. 1.
    Define a nonconformity score function $s(x,y)\in\mathbb{R}$ that quantifies the disagreement between the model’s prediction for input $x$ and a ground-truth label $y$ (e.g. the absolute difference between the predicted and expected DSC score). Higher scores indicate poorer fit or higher error.
  2. 2.
    Compute $\hat{q}$ as the $\lceil(1-\alpha)(1+1/n)\rceil$-th empirical quantile of the calibration scores $\{s(x_{1},y_{1}),\ldots,s(x_{n},y_{n})\}$. For example, for a desired coverage of 90% ($\alpha=10\%$), $\hat{q}$ will be the nonconformity score under which 90% of the nonconformity scores computed in the calibration set fall.
  3. 3.
    Construct the prediction set for $x_{n+1}$ as:
    $C_{\alpha}(x_{n+1})=\{y:s(x_{n+1},y)\leq\hat{q}\}$.
    This set (or interval in case of regression problems) will contain all candidate outputs for $x_{n+1}$ that have a nonconformity score less than or equal to the threshold.
This procedure gives the following coverage guarantee:
Theorem 1
Suppose that $\{(X_{i},Y_{i})\}_{i=1}^{n}$ and $(X_{n+1},Y_{n+1})$ are i.i.d. Let $\hat{q}$ and $C_{\alpha}(X_{n+1})$ be as defined above. Then the following holds:
$\mathbb{P}(Y_{n+1}\in C_{\alpha}(X_{n+1}))\geq 1-\alpha.$ (3)
The proof of this result traces back to [7]. Although the coverage guarantee holds regardless of the score function, the usefulness of the prediction sets is determined by how well it captures the model’s uncertainty. In practice, conformal prediction methods are primarily evaluated in terms of validity (i.e., achieving the desired coverage level) and efficiency (i.e., producing prediction sets or intervals that are as tight as possible).
In regression problems, a widely adopted nonconformity score is the residual $s(x,y)=|y-\hat{f}(x)|$, where $\hat{f}(x)$ denotes the model’s prediction for input $x$ [7]. The resulting conformal prediction interval takes the form $C_{\alpha}(x_{n+1})=\hat{f}(x_{n+1})\pm\hat{q}$. Note that this leads to prediction intervals of equal width for all test inputs, regardless of their location in the input space. To better account for heteroscedasticity, i.e. when the noise variance varies across the input domain, a locally weighted nonconformity score has been proposed: $s(x,y)=\frac{|y-\hat{f}(x)|}{\hat{\sigma}(x)}$, where $\hat{\sigma}(x)$ is the estimated standard error of the residual [41].
An alternative and popular approach that naturally accounts for heteroscedasticity is Conformalized Quantile Regression (CQR) [11]. CQR, builds on quantile regression models, which are trained to predict the conditional lower and upper quantiles $\hat{t}_{\alpha/2}(x)$ and $\hat{t}_{1-\alpha/2}(x)$ of the response $Y$ given an input
$X=x$. The nonconformity score is defined as the distance from $y$ to the nearest quantile:
$s(x,y)=\max\left\{\hat{t}_{\alpha/2}(x)-y,\;y-\hat{t}_{1-\alpha/2}(x)\right\}$
and the prediction interval for a new input $x_{n+1}$ is constructed as:
$C_{\alpha}(x_{n+1})=\left[\hat{t}_{\alpha/2}(x_{n+1})-\hat{q},\;\hat{t}_{1-\alpha/2}(x_{n+1})+\hat{q}\right$].
We build on the CQR framework as the basis for our proposed method. This approach maintains the coverage guarantee of CP while often producing tighter, more locally adaptive intervals compared to residual-based methods, particularly in heteroscedastic settings [11].
IV-B Proposed Method
Traditional RCA yields a point estimate of segmentation quality for an individual image, as defined in Eq. (2). However, it does not quantify the uncertainty associated with this prediction. To address this limitation, we extend RCA following a split conformal prediction approach, producing prediction intervals with statistical guarantees.
Let $\mathcal{C}=\{(I_{j}^{C},S_{I_{j}^{C}},S_{I_{j}^{C}}^{GT})\}_{j=1}^{n}$ be a calibration set of images with both predicted ($S_{I_{j}^{C}}$) and ground-truth ($S_{I_{j}^{C}}^{GT}$) segmentations. For each calibration sample $(I_{j}^{C},S_{I_{j}^{C}})$, we apply the RCA procedure: we train a reverse segmenter $f_{I_{j}^{C},S_{I_{j}}^{C}}$ using $S_{I_{j}^{C}}$ as pseudo-ground truth, evaluate it on the reference database $D$, and obtain the corresponding RCA scores $\mathcal{Y}_{j}^{C}$ following Eq. (1). We also compute the true quality score $y(S_{I_{j}^{C}})=\rho(S_{I_{j}^{C}},S_{I_{j}^{C}}^{GT})$ of each sample in the calibration set.
To construct the prediction intervals, we follow an approach inspired by CQR but differing in a key aspect: instead of training a quantile regression model, we compute empirical quantiles directly from the RCA score distribution $\mathcal{Y}$. This allows us to naturally capture the uncertainty inherent in the RCA procedure while maintaining computational efficiency.
From each RCA score distribution $\mathcal{Y}_{j}^{C}$, we compute empirical quantiles at asymmetric levels that are chosen to reflect RCA’s characteristic behavior. Specifically, the lower quantile $q_{\ell}(S_{I_{j}^{C}})=\text{Quantile}(\mathcal{Y}_{j}^{C},p_{\ell})$ is selected near the center of the distribution, while the upper quantile $q_{h}(S_{I_{j}^{C}})=\text{Quantile}(\mathcal{Y}_{j}^{C},p_{h})$ is chosen closer to the maximum. This asymmetry captures the observation that RCA tends to underestimate segmentation quality, making higher scores more informative [5].
We then define nonconformity scores that measure how far the true quality deviates from the RCA-based interval bounds:
$s_{j}=\max\{q_{\ell}(S_{I_{j}^{C}})-y(S_{I_{j}^{C}});\ y(S_{I_{j}^{C}})-q_{h}(S_{I_{j}^{C}})\}.$ (4)
Based on the calibration nonconformity scores, $\{s_{j}\}_{j=1}^{n}$, the “CP threshold”, $\hat{q}$, is computed as the $\lceil(1-\alpha)(1+1/n)\rceil$-th smallest value among $\{s_{j}\}_{j=1}^{n}$ (see Step 2 in Section IV-A). For a new test pair $(I,S_{I})$, we compute its RCA scores $\mathcal{Y}$ and compute the corresponding empirical quantiles $q_{\ell}(S_{I})$ and $q_{h}(S_{I})$. The conformalized prediction interval is then constructed as:
$C(S_{I})=\left[q_{\ell}(S_{I})-\hat{q},q_{h}(S_{I})+\hat{q}\right$]. (5)
Assuming exchangeability between the test sample and calibration set, the constructed interval provides the marginal coverage guarantee stated in Theorem 3, as shown in [11]. In our experiments, $\rho$ corresponds to the Dice Similarity Coefficient (DSC), which takes values in the $[0,1]$ range. Accordingly, we clip the predicted interval $C(S_{I})$ to this range to ensure plausible quality estimates. Note that the procedure can be applied with any reverse segmenter and can thus be readily combined with the proposed in-context approach.
Figure 3: Scatter plots comparing predicted vs real DSC on 3D-IRCAdB (Left) and JSRT (Right) for different reference datasets of size $k$ using UniverSeg as the reverse classifier. Top: results using randomly selected reference samples. Bottom: results following retrieval-augmentation, where the $k$ most similar images are selected based on cosine similarity in the DINOv2 embedding space. In both cases retrieval-augmentation achieves better performance with fewer samples, as shown by higher correlation and lower MAE (bold indicates the best value).
Figure 5: (a) Distribution of prediction interval sizes (the lower the better) and (b) empirical coverage rates (the larger the better) for conformal prediction methods across all evaluated datasets. Results compare our Conformal-RCA variants using different reverse segmenters (Atlas-based, UniverSeg, SAM 2) against the baseline method PHiSeg from [26], as well as our proposed variant PHiSeg (CQR), which follows the empirical CQR approach. Target coverage level is set to 90% ($\alpha=0.1$).
V Experiments and results
V-A Datasets & Computing Resources
We evaluate the proposed Conformal and In-Context RCA variants across diverse datasets listed in Fig. 2 including SCD [42], HC18 [43], PSFHS [44], PH2 [45], ISIC 2018 [46], NuCLS [47], JSRT [48], 3D-IRCAdB [49], WBC-CV and WBC-JTSC [50]. Unlike previous studies that explored RCA focusing on cardiac MRI [19, 20] or multi-organ segmentation on whole-body MRI [5] only, we broaden our analysis to a wider range of medical imaging challenges. To maintain consistency, we convert all images to grayscale and resize them to 256x256 pixels. For 3D images, we extract slices, keeping those where the region of interest (ROI) is present.
All the experiments that will be described here were conducted on an AMD Ryzen 7 4800H CPU (8 cores) and an NVIDIA GTX 1660 Ti GPU (6GB of VRAM).
V-B Evaluation of In-Context RCA